Show that the functions, defined by f:R->(1,infinity) ->R f(x)=3^2x+1, g(x)=log(x-1) are inverse of one another.
1
Expert's answer
2021-09-21T11:45:18-0400

Solution:

"f(x)=3^{2x}+1"

"Domain: (-\infin, \infin)"

"Range: (1, \infin)"

Replace "f(x)" by "y"



"y=3^{2x}+1"

Interchange "x" and "y"



"x=3^{2y}+1"

Solve for "y"



"3^{2y}=x-1""y=\dfrac{1}{2}\log_3(x-1)"

Replace "y" by "f^{-1}(x)"



"f^{-1}(x)=\dfrac{1}{2}\log_3(x-1)"

"Domain: (1, \infin)"

"Range: (-\infin, \infin)"

The functions "f(x)=3^{2x}+1" and "f^{-1}(x)=\dfrac{1}{2}\log_3(x-1)" are inverse of one another.

Hence, proved.